OpenStudy (anonymous):

The half life of phospherous 32 is about 14 days. There are 6.6 grMA INITIALLY. Express the amount of phospherous 32 remaining as a function of time (t). We have to use the equation y=ae^-kt

OpenStudy (anonymous):

\[A(t)=6.6\times \left(\frac{1}{2}\right)^{\frac{t}{14}}\] will work, without much effort

OpenStudy (anonymous):

when i would solve the equation i would get a very small numbe. Ok i had that set up wrong

OpenStudy (anonymous):

your teacher won't like it though, they want you to use base \(e\) and solve for \(k\) first

OpenStudy (anonymous):

we can do it that way too if you like, i am sure that is what you are supposed to do

OpenStudy (anonymous):

So how would i exactly get that t/14 by itself?

OpenStudy (anonymous):

you know the half life is \(14\) days you the first step is to solve \[e^{14t}=\frac{1}{2}\] for \(t\)

OpenStudy (anonymous):

oh i have confused you my initial answer \(A(t)=6.6\times \left(\frac{1}{2}\right)^{\frac{t}{14}}\) is also my final answer it is an alternative to \(A_0e^{kt}\)

OpenStudy (anonymous):

ok so i just have to ln both sides to cancel e

OpenStudy (anonymous):

lets do it the way your teacher wants

OpenStudy (anonymous):

solve \[e^{14k}=\frac{1}{2}\] because you know it 14 days, i.e. when \(t=14\) you have half of what you started with

OpenStudy (anonymous):

solve quickly \[e^{14k}=.5\] \[14k=\ln(.5)\] \[k=\frac{\ln(.5)}{14}\]

OpenStudy (anonymous):

Ok i was multiplying by taht 14

OpenStudy (anonymous):

you can easily generalize this now we know \(k=-0.0495\) rounded

OpenStudy (anonymous):

and your final answer is \[A(t)=6.6e^{-.0495t}\]

OpenStudy (anonymous):

ok so the i put14 in for t and taht answer in for k and solve the equation?

OpenStudy (anonymous):

hold on, i sense you are maybe confused, so lets go slow

OpenStudy (anonymous):

ya im not really sure how to solv t

OpenStudy (anonymous):

this is the question "Express the amount of phospherous 32 remaining as a function of time (t). We have to use the equation \(y=ae^{-kt}\)"

OpenStudy (anonymous):

you are not asked to solve for \(t\) you are asked to find the function that gives the amount at time \(t\) the function is a function of \(t\) so you need a \(t\) in the answer

OpenStudy (anonymous):

your job was to take the given information and solve for \(\huge k\)

OpenStudy (anonymous):

one you have it, you are done, you replace \(k\) by your answer in the function \[y=ae^{kt}\] and you are finished

OpenStudy (anonymous):

you solve for \(t\) when you are asked something like "how long before..." or "how old is..." that is when you want the time

OpenStudy (anonymous):

ok i i got that i already know t k is my unknown. Thanks

OpenStudy (anonymous):

well you do not "know" \(t\) as \(t\) is variable, but yes, you need to solve for \(k\) to get the function i hope the steps to find it were clear ignore my first answer, i think it confused the issue

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

yw